# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

### v4

@@ 4.1.1. Input Language and Modality
Sankara Eye Hospital in Bangalore (in the state of Karnataka), referred to as the Silicon Valley of India, caters to patients from various linguistic, educational, and technical backgrounds, including those from the Information Technology sector and neighboring states. Analysis of a 2011 census highlights Bangalore as one of India’s most linguistically diverse cities (Office of the Registrar General & Census Commissioner, 2011). To accommodate this varying patient demography, CataractBot is designed to support five languages: English, Hindi (the local language of nine states in India), Kannada (the local language of Karnataka), and Tamil and Telugu (the local languages of two neighboring states of Karnataka). As the expert may not be proficient in all five languages, their interactions with the bot are exclusively in English. Upon onboarding, the language preference of the patient and their attendant is collected through an online form, and CataractBot initiates the conversation by sending a set of ‘welcome messages’ in the chosen language. Among these messages is an option allowing users to modify their preferred language in future. To engage with users with different levels of literacy, including literate, semi-literate, and illiterate users, CataractBot supports both text and speech inputs. For every voice message, CataractBot sends both a text response and a spoken version of the text as an audio message.
Recognizing the challenge users face in initiating a conversation, often referred to as ‘Blank Page Syndrome’ (Moden, 2022; YM et al., 2023), CataractBot provides a set of three frequently asked questions along with the welcome message. Additionally, every response includes a “What to do next?” prompt, offering three related questions based on the preceding query (Figure 2F). Users can select one of the suggested questions or pose a different query to continue the conversation. This feature lowers the barrier to entry (as formulating questions has been found to be non-trivial for low-literate individuals (Tseng et al., 2023)), enhances the natural flow of conversation, and saves time.
@@ 4.1.2. Response Generation, Verification and Icons
The response includes a question mark icon
as a ‘reaction’ indicating the unverified status (Figure 2E).
@@
For medical questions, the patient’s operating doctor receives a message comprising of the question asked, the bot’s response, and patient demographics (Figure 1A).
@@
The doctor is prompted with the question, “Is the answer accurate and complete?” offering three response options: Yes, No, and Send to Patient Coordinator (Figure 1A).
@@
, confirming verification (Figure 1F). Additionally, the bot notifies the information seeker that the answer has been verified, tagging that particular response (Figure 1G).
@@
The system automatically combines the bot’s initial answer with the doctor-provided correction to frame a new response, which is delivered to the user with a green tick icon
@@
Similar to the doctors’ workflow, patient coordinators verify and provide corrections to the bot’s responses for logistical questions, and have the option to ‘Send to Doctor’ if a medical question is misclassified.
@@
Delayed responses might lead to patients making repeated calls to the hospital or resorting to online sources, failing to effectively address their information needs.
@@ 4.1.3. Escalation Mechanism
(Figure 2E) and answered questions with a green tick
@@
If an answer is not verified by the operating doctor within three hours, it is automatically sent to the designated escalation doctor.
The operating doctor receives a notification about the escalation.
@@
This additional reminder was added based on the doctors’ feedback.
This workflow is mirrored for the patient coordinator and escalation patient coordinator.
@@ 4.1.4. Knowledge Base Update Process
To minimize experts’ labor, we use the expert-provided edits to update the knowledge base, which increases the likelihood of ‘Yes’ responses from experts for similar questions in the future. However, certain responses, such as those specific to individual patients (e.g., “When to reach the hospital for my surgery?”) must not be added to the knowledge base as they are not generalizable. We appointed a senior cataract surgeon as the ‘knowledge base expert’, who received a spreadsheet via email at 8pm daily. Their task involves reviewing each row, which consists of question-answer pairs, and determining whether the information should be added to the knowledge base (by responding with a ‘Yes’/‘No’ in the ‘Should Update Knowledge Base?’ column), and if so, modifying as needed the ‘Final Answer for Knowledge Base’ column, containing the bot’s updated answer based on the expert’s correction. At 3 am daily, the system extracts ‘Question’ and ‘Final Answer for Knowledge Base’ data from all rows marked ‘Yes’, and append these to the knowledge base in the ‘expert-FAQ’ document. Additionally, we prompt the LLM to prioritize this ‘expert-FAQ’ document within the knowledge base. This ensures that CataractBot improves with these updates, gradually enhancing its accuracy over time.
@@ 4.2. Implementation Details
We opted for GPT-4, the leading LLM at the time of our system development in May 2023 (OpenAI et al., 2023). The custom knowledge base comprises of various cataract surgery and hospital-specific documents, including Consent Form, SOP (Standard Operating Procedure), FAQs, Pre- and Post-Operative Guidelines, and Hospital Information. These documents are ingested as data chunks into a Chroma vector database, using the OpenAI model ‘text-embedding-ada-002’ to generate embeddings. Upon receiving a question, GPT-4 classifies the query type. For a medical/logistical question, CataractBot employs a retrieval-augmented generation (Lewis et al., 2020) approach. This involves performing a vector search on the knowledge base to extract the three most relevant data chunks related to the query. GPT-4 is then prompted (full prompt available in Appendix A) to generate an answer for the query from these data chunks. If the answer is not present, the bot responds with an “I don’t know” message.
@@ 5. Study Design
To evaluate the effectiveness of this LLM-powered experts-in-the-loop chatbot in addressing the informational needs of patients undergoing cataract surgery, we conducted a mixed-method user study during Nov 2023-Jan 2024 at our collaborators’ institute, Sankara Eye Hospital. This is one of the leading tertiary eye care and teaching institutions in Bangalore, India. It typically attends to more than 500 patients every day, and conducts over 50 cataract surgeries daily. This study was approved by both the Scientific and Ethics Committees of Sankara Eye Hospital.. None of the study participants—patients, attendants, doctors, or coordinators—received any financial incentives for their participation, in accordance with hospital norms.
@@ 5.1. Procedure for Patients and Attendants
As per the hospital’s protocol, once a patient opts for cataract surgery based on their doctor’s recommendation, the patient and their attendant needs to meet a patient coordinator. The coordinator helps with general guidance on pre- and post-operative measures, scheduling of the surgery date, and logistical arrangements. Toward the end of this interaction, the coordinator was tasked to assess the patient’s eligibility for the study based on specific criteria: (a) age of 18 or older, (b) fluent in one of the five languages supported by CataractBot, (c) scheduled for cataract surgery within a week, (d) no history of cataract surgery in the last 6 months (as recent patients would likely have minimal informational needs), and (e) having one of the three participating operating doctors as their surgeon. If these criteria were met, the patient and their attendant were directed by the coordinator to meet a researcher (the first author) stationed at the hospital.
@@
The researcher introduced them to CataractBot and outlined the study’s protocol which involved using the bot for around two weeks, engaging in an interview pre and post-surgery, and completing a chatbot usability form. If they agreed to participate, participants were requested to sign a consent form, after which the researcher filled a web-based onboarding form. This form included details such as the patient and attendant’s phone numbers linked to WhatsApp, preferred languages, consulting doctor and coordinator, surgery date, and basic demographic information (age, gender, and education). Upon form submission, participants received ‘welcome messages’ from CataractBot. The researcher instructed them to ask a trial question, either by choosing from suggested questions or by typing/speaking in their preferred language. Upon receiving a response, participants were briefed on the icons and expert verification system. Throughout this process, the researcher encouraged them to ask any questions regarding the bot’s usage. After onboarding, participants received two daily reminder messages (at 7:30 am and 4 pm) prompting them to “Feel free to ask any cataract surgery related questions” until one week post surgery, when their access was automatically revoked.
@@
We conducted two semi-structured interviews with both the patient and their attendant: one on the surgery day and another a week post-surgery. These specific days were chosen because they coincided with the patient’s required hospital visits. The interviews explored their overall experience, specific features they liked or disliked, suggestions for improvement, and questions regarding the trustworthiness, timeliness, and accuracy of responses. Both interviews followed a similar structure, differing only in how the bot supported the participant before versus after surgery. After each interview, participants were requested to fill out a chatbot usability form. They rated their chatbot experience through eight questions, using a five-point Likert scale, covering aspects like ease of use, understandability, and timeliness (Appendix A.2). These questions were adapted from metrics outlined in prior usability evaluations of healthcare bots; previous work used scales that included aspects not relevant to our study, such as the chatbot’s visual design (Abd-Alrazaq et al., 2020; Holmes et al., 2023). Patient-attendant pairs were interviewed together but were asked to fill separate usability forms.
@@ 5.3. Participants
A total of 31 patient-attendant pairs (Table 2 in Appendix A.3), consisting of 19 patients and 30 attendants, took part in our study. Due to the predominantly elderly demographic undergoing cataract surgery, 12 patients lacked access to WhatsApp/smartphones and were unable to participate but their attendants used the bot. One patient was not accompanied by an attendant. The average age of the 19 participating patients (3 females) was 58.88.01 years. Although females comprised 32.7% of our study participants, only 15.8% of the patients were female. This aligns with prior work showing a gender disparity in accessing cataract surgery in India (Prasad et al., 2020; Ye et al., 2020). Patients had diverse preferred languages (5 English, 5 Hindi, 5 Kannada, 3 Tamil, and 1 Telugu) and education levels (6 Grade 10, 4 Grade 12, 6 Bachelors, and 3 Masters). In contrast, the attendants were younger, predominantly fluent in English, and well-educated. The average age of the 30 participating attendants (11 female) was 37.210.3 years. A majority (23) preferred English, with 2 each favoring Hindi, Kannada, and Tamil, and 1 preferring Telugu. Education levels were mostly high, with 10 having Masters and 13 having Bachelors degrees. Among the 31 patient-attendant pairs, 22 participated in the surgery-day interview, and 10 took part in the one-week post-surgery interview. Various factors influenced participation, including time constraints (resulting in a few declined interviews) and data reaching saturation. Additionally, four doctors and two patient coordinators participated in the study (Table 3 in Appendix A.3). The roles of escalation doctor and knowledge base expert were performed by the same doctor. Note: there was a partial overlap between participants in this study and the formative study, with two doctors and one coordinator participating in both.
@@ 5.4. Data collection and analysis
We performed a mixed-method analysis to systematically analyze the collected data comprising of usage logs, interview transcripts, and usability survey responses. CataractBot interaction logs and usability survey data were quantitatively analyzed. Descriptive statistics and statistical tests (such as t-tests and ANOVAs) were used to analyze the count and type of questions asked, verification styles and response times of doctors and coordinators, edits performed by the knowledge base update expert, and responses to the usability survey. Note: For ANOVAs, the sphericity assumption was tested using Mauchly’s test, and in case of sphericity violations, the Greenhouse-Geisser correction was applied.
The first author translated and transcribed all interviews (totalling 11.5 hours) into English soon after they were conducted. Throughout the study, all authors regularly engaged in discussions to review observations and interviews. To analyze interview transcripts, we employed an inductive coding approach, with the first author open-coding (Williams and Moser, 2019) all interviews line-by-line. Subsequently, two authors collaboratively reviewed these codes to identify interesting themes in the data (similar to (Okolo et al., 2021)). We iterated on these themes to distill higher-level themes (like “Usage”, “Features”, and “Privacy and Trust”) that we present in our findings below. Note: While there were a total of 49 patients and attendants using CataractBot, at times, we treat them as 31 participants (patient-attendant pairs). This choice stems from instances highlighted during interviews, where patients instructed attendants to query the bot using their phones. Additionally, both were interviewed together. Only when comparing their distinct roles, we analyzed their usage log data separately.
@@ 5.5. Positionality
All authors are of Indian origin, and currently reside in Bangalore, India. Among them, two identify as female and the rest as male. One author is a practising ophthalmologist, performing over 10 cataract surgeries weekly, and another author is part of the Quality Control and Patient Safety team at the same hospital. Both were involved in designing the study, shaping research questions, providing feedback on initial bot versions, curating the knowledge base, and ensuring that any risk of harm to the patients was mitigated. The ophthalmologist also served as the knowledge base expert. Two authors specialize in HCI and design, three possess healthcare research experience, and four have expertise in software development. We approached this research drawing on our individual experiences and learning from working at the intersection of computing and healthcare in the Indian context. Our interest in making healthcare more accessible to the masses has informed the design of CataractBot, guided our study design, and shaped our data analysis.
@@ 6. Findings
Overall, 31 patient-attendant pairs sent 343 messages to CataractBot, with an average of 11.068.44 messages per pair. Out of the 343 messages, 225 were medical questions, 87 logistical, 27 small talk, and 4 others. Doctors directly approved (said ‘Yes’ to) 69.8% (157) of the bot-generated answers to medical questions while patient coordinators directly approved 58.6% (51) of logistical answers. In the usability survey, 77.3% of participants found CataractBot’s responses to be useful, appropriate, and informative, while 91.0% said they would be willing to use it in the future. Below, in our findings, we delve into how participants used CataractBot (answering RQ1 and RQ3), examine the impact of novel features introduced in CataractBot (answering RQ2), and explore privacy and trust implications of using an LLM-powered chatbot (answering RQ1).
6.1. CataractBot Usage

### v5

@@ 4. CataractBot System Design
| Stakeholder | Role | ||||||
| Information seeker | Patient |
| Attendant |
| Expert | Doctor |
| Patient coordinator |
@@ 4.1.1. Input Language and Modality
Sankara Eye Hospital in Bangalore (in the state of Karnataka), referred to as the Silicon Valley of India, caters to patients from various linguistic, educational, and technical backgrounds, including those from the Information Technology sector and neighboring states. Analysis of a 2011 census highlights Bangalore as one of India’s most linguistically diverse cities (Office of the Registrar General & Census Commissioner, 2011). To accommodate this varying patient demography, CataractBot is designed to support five languages: English, Hindi (the local language of nine states in India), Kannada (the local language of Karnataka), and Tamil and Telugu (the local languages of two neighboring states of Karnataka). As the expert may not be proficient in all five languages, their interactions with the bot are exclusively in English. Upon onboarding, the language preference of the patient and their attendant is collected through an online form, and CataractBot initiates the conversation by sending a set of ‘welcome messages’ in the chosen language. Among these messages is an option allowing users to modify their preferred language in future. To engage with users with different levels of literacyCataractBot supports both text and speech inputs. For every voice message, CataractBot sends both a text response and a spoken version of the text as an audio message.
Recognizing the challenge users face in initiating a conversation, (‘Blank Page Syndrome’ (Moden, 2022; YM et al., 2023)), CataractBot provides a set of three frequently asked questions along with the welcome message. Additionally, every response includes a “What to do next?” prompt, offering three related questions based on the preceding query (Figure 4F in Appendix A.1.2). Users can select one of the suggested questions or pose a different query to continue the conversation. This feature lowers the barrier to entry (as formulating questions has been found to be non-trivial for low-literate individuals (Tseng et al., 2023)), enhances the natural flow of conversation, and saves time.
@@ 4.1.2. Response Generation, Verification and Icons
This custom knowledge base comprises of various cataract surgery and hospital-specific documents curated by hospital staff (doctors, patient coordinators, and members of Quality Control and the Patient Safety team).
These documents, comprising approximately 30 pages in total, include the Consent Form, Standard Operating Procedures, FAQs, Pre- and Post-Operative Guidelines, and Hospital Information.
Some documents, like Post-Operative Guidelines, were included in patients’ discharge summaries.
Others, such as SOPs and FAQs, existed in the hospital ecosystem but were not directly accessible to patients.
The generated response includes a question mark icon
as a ‘reaction’ indicating the unverified status (Figure 4E in Appendix A.1.2).
@@
For medical questions, the patient’s operating doctor (Table 1) receives a message comprising of the question asked, the bot’s response, and patient demographics (Figure 1A).
@@
The doctor is prompted with the question, “Is the answer accurate and complete?” offering three response options: “Yes”, “No”, and “Send to Patient Coordinator”.
@@
, confirming verification (Figure 1F).
Additionally, the bot notifies the information seeker that the answer has been verified, tagging that particular response (Figure 1G).
@@
The system combines the bot’s initial answer with the doctor-provided correction to frame a new response, which is delivered to the user with a green tick icon
@@
Similar to the doctors’ workflow, patient coordinators verify and provide corrections to the bot’s responses for logistical questions.
@@
Delayed responses might lead to patients making repeated calls to the hospital or resorting to online sources, failing to meet their information needs.
@@ 4.1.3. Escalation Mechanism
(Figure 4E in Appendix A.1.2) and answered questions with a green tick
@@
If an answer is not verified by the operating doctor within three hours, it is automatically sent to the designated escalation doctor (Table 1).
@@
This reminder was added based on the doctors’ feedback.
This workflow is mirrored for the operating and escalation patient coordinators.
@@ 4.1.4. Knowledge Base Update Process
To minimize experts’ labor, we use expert-provided edits to update the knowledge base, which increases the likelihood of ‘Yes’ responses from experts to similar questions in the future. However, certain responses, such as those specific to individual patients (e.g., “When to reach the hospital for my surgery?”) must not be added to the knowledge base as they are not generalizable. We appointed a senior cataract surgeon as the ‘knowledge base expert’ (Table 1), who received a spreadsheet via email at 8pm daily. Their task involves reviewing each row, which consists of question-answer pairs, and determining whether the information should be added to the knowledge base (by responding with a ‘Yes’/‘No’ in the ‘Should Update Knowledge Base?’ column), and if so, modifying as needed the ‘Final Answer for Knowledge Base’ column, containing the bot’s updated answer based on the expert’s correction. At 3 am daily, the system extracts ‘Question’ and ‘Final Answer for Knowledge Base’ data from all rows marked ‘Yes’, and append these to the knowledge base in the ‘expert-FAQ’ document. Additionally, we prompt the LLM to prioritize this ‘expert-FAQ’ document within the knowledge base. This ensures that CataractBot improves with these updates, gradually enhancing its accuracy over time.
@@ 4.2. Implementation Details
We opted for GPT-4, the leading LLM at the time of our system development in May 2023 (OpenAI et al., 2023). The documents for the custom knowledge base are ingested as data chunks into a Chroma vector database, using the OpenAI model ‘text-embedding-ada-002’ to generate embeddings. Upon receiving a question, GPT-4 classifies the query type. For a medical/logistical question, CataractBot employs a retrieval-augmented generation (Lewis et al., 2020) approach. This involves performing a vector search on the knowledge base to extract the three most relevant data chunks. GPT-4 is then prompted (full prompt available in Appendix A.1.1) to generate an answer for the query from these data chunks. If the answer is not present, the bot responds with an “I don’t know” message.
@@ 5. Study Design
To evaluate the effectiveness of this LLM-powered experts-in-the-loop chatbot in addressing the informational needs of patients undergoing cataract surgery, we conducted a mixed-method user study during Nov 2023-Jan 2024 at our collaborators’ institute, Sankara Eye Hospital. This is one of the leading tertiary eye care and teaching institutions in Bangalore, India. It typically attends to more than 500 patients every day, and conducts over 50 cataract surgeries daily. This study was approved by both the Scientific and Ethics Committees of Sankara Eye Hospital. None of the study participants—patients, attendants, doctors, or coordinators—received any financial incentives for their participation, in accordance with hospital norms.
@@ 5.1. Procedure for Patients and Attendants
The operating patient coordinator was tasked to assess the patient’s eligibility for the study based on specific criteria: (a) age of 18 or older, (b) fluent in one of the five languages supported by CataractBot, (c) scheduled for cataract surgery within a week, (d) no history of cataract surgery in the last 6 months (as recent patients would likely have minimal informational needs), and (e) having one of the three participating operating doctors as their surgeon. If these criteria were met, the patient and their attendant were directed by the coordinator to meet a researcher (the first author) stationed at the hospital.
@@
The researcher introduced them to CataractBot and outlined the study’s protocol which involved using the bot for around two weeks and engaging in an interview pre- and post-surgery. If they agreed to participate, participants were requested to sign a consent form, after which the researcher filled a web-based onboarding form. This form included details such as the patient and attendant’s phone numbers linked to WhatsApp, preferred languages, consulting doctor and coordinator, surgery date, and basic demographic information (age, gender, and education). Upon form submission, participants received ‘welcome messages’ from CataractBot. The researcher instructed them to ask a trial question, either by choosing from suggested questions or by typing/speaking in their preferred language. Upon receiving a response, participants were briefed on the icons and expert verification system. Throughout this process, the researcher encouraged them to ask any questions regarding the bot’s usage. After onboarding, participants received two daily reminder messages (at 7:30 am and 4 pm) prompting them to “ask any cataract surgery related questions” until one week post surgery, when their access was revoked.
@@
We conducted two semi-structured interviews: one on the surgery day and another a week post-surgery. These specific days were chosen because they coincided with the patient’s required hospital visits. The interviews explored their overall experience, specific features they liked or disliked, suggestions for improvement, and questions regarding the trustworthiness, timeliness, and accuracy of responses. Both interviews followed a similar structure, differing only in how the bot supported the participant before versus after surgery. After each interview, participants were requested to fill out a chatbot usability form. Those who agreed, rated their chatbot experience through eight questions, using a five-point Likert scale, covering aspects like ease of use, understandability, and timeliness (Appendix A.3). These questions were adapted from metrics outlined in prior usability evaluations of healthcare bots (Abd-Alrazaq et al., 2020; Holmes et al., 2023). Patient-attendant pairs were interviewed together but filled separate usability forms.
@@ 5.3. Participants
A total of 31 patient-attendant pairs (Table 3 in Appendix A.2), consisting of 19 patients and 30 attendants, took part in our study. Due to the predominantly elderly demographic undergoing cataract surgery, 12 patients lacked access to WhatsApp/smartphones and were unable to participate but their attendants used the bot. One patient was not accompanied by an attendant. The average age of the 19 patients was 58.88.01 years. Although females comprised 18 (32.7%) of all participants, only 3 (15.8%) were patients. This aligns with prior work showing a gender disparity in accessing cataract surgery in India (Prasad et al., 2020; Ye et al., 2020). Patients had diverse preferred languages (5 English, 5 Hindi, 5 Kannada, 3 Tamil, and 1 Telugu) and education levels (6 Grade 10, 4 Grade 12, 6 Bachelors, and 3 Masters). In contrast, the attendants were younger, predominantly fluent in English, and well-educated. The average age of the 30 attendants (11 female) was 37.210.3 years. A majority (23) preferred English, with 2 each favoring Hindi, Kannada, and Tamil, and 1 preferring Telugu. Education levels were mostly high, with 10 having Masters and 13 having Bachelors degrees. Among the 31 patient-attendant pairs, 22 participated in the surgery-day interview, and 10 took part in the one-week post-surgery interview. Various factors influenced participation, including time constraints (resulting in a few declined interviews) and data reaching saturation. Additionally, four doctors and two patient coordinators participated in the study (Table 4 in Appendix A.2). The roles of escalation doctor and knowledge base expert were performed by the same doctor. Note: there was a partial overlap with two doctors and one coordinator participating in both this study and the formative study.
@@ 5.4. Data collection and analysis
We performed a mixed-method analysis to systematically analyze the collected data comprising of usage logs, interview transcripts, and usability survey responses. CataractBot interaction logs and usability survey data were quantitatively analyzed. Descriptive statistics and statistical tests (such as t-tests and ANOVAs) were used to analyze the count and type of questions asked, verification styles and response times of doctors and coordinators, edits performed by the knowledge base update expert, and responses to the usability survey. Note: For ANOVAs, the sphericity assumption was tested using Mauchly’s test, and in case of sphericity violations, the Greenhouse-Geisser correction was applied. The first author translated and transcribed all interviews (totalling 11.5 hours) into English soon after they were conducted. Throughout the study, all authors regularly engaged in discussions to review observations and interviews. For interview transcripts, we conducted an inductive thematic analysis (Braun and Clarke, 2006) approach, with the first author open-coding all interviews line-by-line. Subsequently, two authors collaboratively reviewed these codes to identify interesting themes in the data (similar to (Okolo et al., 2021)). We iterated on these themes to distill higher-level themes (like “Usage” and “Features”), that we present in our findings below. Note: While there were a total of 49 patients and attendants using CataractBot, at times, we treat them as 31 participants (patient-attendant pairs). This choice stems from instances highlighted during interviews, where patients instructed attendants to query the bot using either of their phones. Additionally, both were interviewed together. Only when comparing their distinct roles, we analyze their usage log data separately.
@@ 5.5. Positionality
All authors are of Indian origin, and currently reside in Bangalore, India. One author is a practising ophthalmologist, performing over 10 cataract surgeries weekly, and another author is part of the Quality Control and Patient Safety team at the same hospital. Both were involved in designing the study, shaping research questions, providing feedback on initial bot versions, curating the knowledge base, and ensuring that any risk of harm to the patients was mitigated. The ophthalmologist also served as the knowledge base expert. Two authors specialize in HCI and design, three possess healthcare research experience, and four have expertise in software development. We approached this research drawing on our individual experiences and learning from working at the intersection of computing and healthcare in India. Our interest in making healthcare more accessible to the masses has informed the design of CataractBot and guided our study design.
@@ 6. Findings
Overall, 31 patient-attendant pairs sent 343 messages to CataractBot (11.068.44 messages/pair). 225 messages were medical questions, 87 logistical, 27 small talk, and 4 others. A researcher manually classified all questions into 12 categories, identified in a bottom-up manner (similar to (Li et al., 2023; Clarke et al., 2016b)). Several questions fell into multiple categories. Figure 3A illustrates the relative significance of each category before, on, and after the day of surgery. Before the surgery, questions focused on the procedure (like “anaesthesia”, “lens”, “safety”) and admission/discharge logistics, while post-surgery questions centered on medication and recovery (e.g., “washing hair”, “doing yoga”). Doctors directly approved (said ‘Yes’ to) 69.8% (157) of the bot-generated answers while patient coordinators directly approved 58.6% (51). 17 of the 22 (77.3%) usability survey respondents found CataractBot’s responses to be useful, appropriate, and informative, while 20 (91.0%) indicated willingness to use it in the future.
Below, Section 6.1 presents the reasons behind patients’ and attendants’ engagement with the bot (answering RQ1). Section 6.2 examines the impact of key features of the bot, including the LLM, experts-in-the-loop and support for multiple stakeholders, languages and modalities (RQ2). Finally, Section 6.3 describes the integration of CataractBot into the doctor-patient workflow and discusses issues of personalization and accountability (RQ3). Table 5 in Appendix A.4 presents a summary of key findings across sections.
6.1. Bot’s Role in Addressing Information Needs